# Orthogonal-by-construction augmentation of physics-based input-output models

### v2

@@ Abstract
thanks: This project has received funding from the European Defence Fund programme under grant agreement number No 101103386 and has also been supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7061. This work is also partly funded by the European Union (ERC, COMPLETE, 101075836). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Commission or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ 1 Introduction
In recent years, the increasing complexity of engineering systems and the growing performance demands in control applications have intensified the need for accurate nonlinear models. Discrete-time (DT) input-output (IO) models are one of the most commonly applied structures in system identification as they incorporate a broad spectrum of model classes, ranging from linear time-invariant (LTI) (hespanha_linear_2018) and linear parameter-varying (LPV) (toth_modeling_2010) models to various nonlinear structures (schoukens_nonlinear_2019). First-principle (FP) models can be obtained in DT IO form based on known physical laws and engineering insight. While such physics-based models provide interpretable system descriptions, they often capture only the dominant dynamics, while additional high-complexity effects, such as friction properties or aerodynamic forces, are typically neglected.
@@
As an alternative approach, various data-driven identification methods have been developed for modeling nonlinear systems with an IO model structure (schoukens_nonlinear_2019). In particular, recent advances employing deep artificial neural networks (ANNs) have demonstrated superior modeling accuracy compared to conventional approaches (ljung_deep_2020). However, the practical use of ANN-based black-box models in control-oriented applications, e.g., trajectory planning, remains limited due to their lack of physical interpretability (ljung_perspectives_2010). Furthermore, ANN-based models typically exhibit poor extrapolation capabilities beyond the range of the training data, and substantial learning effort is often spent on rediscovering system behaviors that are already well understood from first-principles knowledge.
@@
To address these challenges, different strategies have been proposed in the literature, starting from (light) grey-box modeling (bohlin_practical_2006) till physics-informed neural networks (PINNs) (raissi_physics-informed_2019) and physics-guided neural networks (PGNNs) (daw_physics-guided_2022). One of the most promising directions of these hybrid approaches is model augmentation (schon_multi-objective_2022; hoekstra_learning-based_2025). The augmentation approach aims at combining FP models, i.e., baseline models, with flexible learning components to achieve faster convergence and better model accuracy compared to black-box learning methods (djeumou_neural_2022). Furthermore, model augmentation produces interpretable models with a clear understanding of how the learning component complements the baseline dynamics.
@@
In this paper, we investigate a widely used model augmentation structure in the literature, namely the additive formulation. This approach connects the physics-based and learning components in parallel and the corresponding parameters of the two components are jointly optimized. However, simultaneously tuning the learning-based and physical parameters results in the two ”subcomponents” competing with each other (bolderman_feedforward_2022). As a result, the ANN can learn relations that could be represented by the FP model, while the baseline parameters can be tuned to physically unrealistic values. This effect undermines physical interpretability of the model estimate and can even compromise the extrapolation capabilities of the resulting model.
@@
This challenge was first addressed in bolderman_feedforward_2022 in the context of PGNN-based feedforward control by introducing an additional regularization term into the cost function, penalizing deviations of the baseline parameters from their nominal values. This addition to the cost function effectively limits deviations of the baseline parameters compared to their initial values, hence preventing them from reaching a physically unrealistic parameter domain. Despite its simplicity, the approach has provided good experimental results (bolderman_physics-guided_2024); moreover, the method can be straightforwardly extended for more complex model augmentation structures (hoekstra_learning-based_2025). Another attractive approach is based on an orthogonal projection-based regularization, introduced in kon_physics-guided_2022, also for feedforward control applications. This approach promotes a specific orthogonality between the baseline and learning components via regularization, penalizing when the ANN learns the already known relations represented by the baseline model. This technique has been adapted and generalized for nonlinear system identification in gyorok_orthogonal_2025; however, since these approaches promote orthogonality via regularization, inherently, there exists a trade-off between model accuracy and the desired complementarity. Finding the appropriate trade-off parameter (i.e., regularization coefficient) may not be intuitive. To overcome this, we propose a direct parametrization with guaranteed orthogonality between the baseline and learning components on a selected data set without requiring any trade-off parameter. We also show that under certain conditions for the dataset used to impose such an orthogonality property, the tuned baseline parameters converge to their physically true values.
@@ 2 Problem statement
where $f^{\mathrm{ANN}}$ here represents a fully connected, feedforward neural network with $\theta_{\mathrm{a}}\in\mathbb{R}^{n_{\theta_{\mathrm{a}}}}$ being the collection of its parameters. Alternatively, $f^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}$ can be replaced by any function approximator without loss of generality. Instead of the additive formulation, many other model augmentation structures can be selected from the literature, e.g., see retzler_learning-based_2024. However, the additive, i.e., parallel, formulation offers a transparent model structure with clear separation between the baseline and learning components (hoekstra_learning-based_2025), hence can be an attractive approach for practical applications.
To achieve the best possible data-fit, while simultaneously acquiring as accurate baseline parameters as possible, an efficient approach is to co-estimate $\theta_{\mathrm{b}}$ and $\theta_{\mathrm{a}}$ parameters, as proposed in bolderman_feedforward_2022. Hence, using a data sequence $\mathcal{D}_{N}=\left\{(x_{i},y_{i})\right\}_{i=0}^{N-1}$ generated by (1), the parameters are estimated by minimizing the prediction error loss function expressed as
@@ 4.1 Recovery of the baseline parameters
If the structure of $\delta$ can be inferred from prior physical knowledge, it is possible to specifically design an input sequence that satisfies Condition 4 (at least with respect to the suspected unmodeled terms). This is a common scenario in practical applications. For example, in aerial robotics, the drag force is known to depend quadratically on the relative velocity of the object (huang_leveraging_2024). However, the associated aerodynamic computations are often complex, and this force term is typically left unmodeled. Incorporating these physics-based insights can, however, guide the design of experiments that fulfil Condition 4.
@@ 4.2 Comparison with the standard parametrization
The error of the estimated baseline parameters in (25) is similar to the one in donati_combining_2025, which also follows intuitively. When the dynamics of (14) are recovered on the training data, the exact baseline parameters are found only if the learning component identifies the unmodeled terms and nothing else. If the ANN learns parts that the baseline model could represent, the true baseline parameters can not be retained.
@@ 4.3 Consistency analysis
In the previous derivations, we assumed that the model training results in such parameters for which the exact relations of the data-generating system are recovered on the training set. Now, we will provide certain conditions under which this assumption is satisfied for $N\to\infty$, i.e., we will show consistency of the estimator following the arguments of ljung_convergence_1978. For that, first we require that the data-generating system (1) is stable.
@@
The identification criterion given by (4) satisfies Condition C1 in ljung_convergence_1978, hence the proof of (ljung_convergence_1978, Lemma 3.1) applies for the considered case. $\blacksquare$
@@
In practice, Condition 11 serves only as a technical condition. After acquiring the estimate of the model parameters $\hat{\theta}$, it can be numerically checked whether the given model realization satisfies this requirement. For a strict guarantee of satisfying Condition 11, a so-called stable-by-design parametrization can be applied. See, e.g., revay_recurrent_2024; kon_unconstrained_2025. Extending these methods for the model augmentation setting is an open question and can be the topic of future research.
@@
For linear time-invariant models, the classical persistency of excitation (PE) condition that is characterized solely on the input trajectory in the open loop case, implies a fundamental informativity requirement, i.e., distinguishability of (equivalence sets of) parameters under the data (Gevers2009). For NARX-type models with linear-in-the-parameters form as (2), the PE condition requires the regressor matrix composed of nonlinear basis functions to be of full column rank. Unlike for linear systems, this condition depends not only on the input signal, but also on the resulting output trajectories. Consequently, PE requirements must be assessed at the level of the nonlinear regressor. Considering the full nonlinear model augmentation structure of (3) further complicates the PE condition. In practice, space-filling and model-based excitation designs (see, e.g., (yuhan_space-filling_2025)) can be applied to select inputs for maximizing data information content over the regressor space.
@@
For the proof of (34), see Lemma 4.1 in ljung_convergence_1978. Note that the applied loss function (4) fulfills Condition (4.4) in ljung_convergence_1978. To prove (35), refer back to Theorem 7. With the conditions of Lemma 15, Assumption 6 is trivially satisfied; hence, (35) holds under Condition 4. $\blacksquare$
@@ 4.4 Covariance of the model parameters
Finally, we show that, due to the orthogonality between the baseline and learning components, the proposed parametrization results in zero covariance between the estimated $\theta_{\mathrm{b}}$ and $\theta_{\mathrm{a}}$ parameters. Under the conditions of the consistency results, the asymptotic distribution of $\hat{\theta}^{N}$ can be expressed w.r.t. to $\lim_{N\rightarrow\infty}\hat{\theta}^{N}=\theta^{\ast}\in\Theta^{\ast}$ as $\sqrt{N}(\hat{\theta}^{N}-\theta^{\ast})\in\mathrm{As}\ \mathcal{N}(0,P_{\theta})$ (ljung_system_1998). Then, under the considered quadratic loss function,
@@
where $\psi_{k}(\theta_{\ast})\in\mathbb{R}^{n_{\mathrm{y}}\times n_{\theta}}$ is the Jacobian matrix [^1] Note that $\psi_{k}$ is defined with a different dimensional notation compared to ljung_system_1998 for practical reasons. $\partial\hat{y}_{k}/\partial\theta$. Based on $\mathcal{D}_{N}$ and parameter estimate $\hat{\theta}^{N}$, $P_{\theta}$ can be estimated as
@@ 5.1 Academic simulation example
where the baseline parameters are $\hat{\theta}_{1}$, and $\hat{\theta}_{2}$ with initial values as $\hat{\theta}_{1}^{0}=0.8$, $\hat{\theta}_{2}^{0}=0.03$. Similar physics-based baseline models have high relevance in system identification theory (schoukens_identification_2017). For the learning component, a simple feedforward ANN with 1 hidden layer and 16 neurons is applied, using the hyperbolic tangent (tanh) activation. The ANN parameters are initialized with the Xavier method (glorot_understanding_2010).
Three distinct datasets are generated for training to highlight the importance of data generation, each with $N=1024$ data points. The first approach applies a white noise input signal with $u_{k}\sim\mathcal{N}(0,0.3^{2})$ for generating 512 samples, then the rest of the data points are acquired by using the first half of the input signal multiplied by -1. This results in a symmetric training data set (denoted by $\mathcal{D}_{N}^{\mathrm{(1)}}$), since the even nonlinearities are orthogonal to odd nonlinearities, according to (17), Condition 4 is satisfied. The second set $\mathcal{D}_{N}^{(2)}$ is gathered by applying 1024 input samples directly generated by the previous distribution. This results in a dataset with, most likely, small asymmetries. However, these asymmetries will disappear, and (17) holds for $N\to\infty$. Finally, $\mathcal{D}_{N}^{(3)}$ is acquired by using an asymmetric distribution of $u_{k}\sim\mathcal{N}(-0.01,0.3^{2})$, which will not satisfy Condition 4 when $N\to\infty$. For testing, a data set is constructed with $N_{\mathrm{test}}=1024$ using the same input distribution as in $\mathcal{D}_{N}^{(2)}$. For better demonstration of the results, the test data is kept noise-free. To minimize (4), we applied the Adam optimizer for 500 epochs, followed by the L-BFGS method for 1000 iterations, using the identification pipeline proposed in bemporad_l-bfgs-b_2025. For benchmarking, we applied the standard additive augmentation, as well as the proposed orthogonal-by-construction method, using 10 different initialization points.
@@ 5.2 Physics-motivated simulation example
Numerical parameter values corresponding to the data-generating system are $m=1$ kg, $k_{\mathrm{s}}=1000~\mathrm{N/m}$, $c_{\mathrm{d}}=100~\mathrm{Ns/m}$, $F_{\mathrm{C}}=0.8~\mathrm{N}$, $F_{\mathrm{S}}=1.2~\mathrm{N}$, $v_{\mathrm{S}}=0.02~\mathrm{m/s}$ with a sampling time of $T_{\mathrm{s}}=0.01$ s. These values result in $\theta_{1}\approx 0.48$, $\theta_{2}\approx-0.48$, $\theta_{3}=\theta_{4}\approx 4.8\cdot 10^{-5}$ (all in appropriate SI unit). Similar MSD interconnections have high importance in mechanical engineering practice, as they are used for dynamic modeling a wide range of components, e.g., suspension of a ground vehicle (prabakar_response_2013). These systems are usually described with linear models, which is convenient for system analysis and control design. Hence, we utilize the following physics-based baseline model:
@@
We apply the same ANN structure, initialization method, and training environment as in Sect. 5.1, but now with 2500 Adam and 500 L-BFGS epochs, respectively. For benchmarking, we apply the standard additive structure, but now complemented with the orthogonal regularization approach (see kon_physics-guided_2022; gyorok_orthogonal_2025), which promotes a similar orthogonality as our method. Moreover, for further benchmarking, we apply another regularization-based technique (denoted as ”baseline regularization” in the remainder), adapted from bolderman_feedforward_2022, which penalizes deviations of the baseline parameters compared to their nominal values. To demonstrate the advantage of our method over regularization-based techniques, we perform a Monte-Carlo study with 5 steps, using various trade-off parameter values (regularization coefficients). As the true baseline model parameters have different magnitudes, now we evaluate the estimation performance by the relative error metric, as $\|(\theta_{\mathrm{b}}^{\ast}-\hat{\theta}_{\mathrm{b}})/\theta_{\mathrm{b}}^{\ast}\|_{2}$. The results are shown in Fig. 5. As shown in the results, the orthogonal regularization–based approach converges toward the proposed method in terms of baseline model parameter estimation error as the trade-off parameter is increased. However, a slight reduction of the overall model performance can be noticed for the regularization-based approach. The baseline regularization method inherently depends on the initial baseline parameter values, since for large values of the trade-off parameter, $\hat{\theta}_{\mathrm{b}}$ remains close to the initial value $\hat{\theta}_{\mathrm{b}}^{0}$. Both approaches demonstrate that, as the regularization coefficient is reduced, the baseline parameters become non-unique and may deviate substantially from the true parameters $\theta_{\mathrm{b}}^{\ast}$. At the same time, lower trade-off parameter values can yield improved model performance, an effect that is particularly visible for the baseline regularization method. In practical system identification scenarios, where only test or validation errors are available during hyperparameter tuning, selecting appropriate regularization coefficients is therefore nontrivial and potentially a complex task. This observation highlights a key advantage of the proposed orthogonal-by-construction method, namely that it requires no hyperparameter tuning.

### v3

@@ Abstract
thanks: This project has been supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7061. This work is also partly funded by the European Union (ERC, COMPLETE, 101075836). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ 1 Introduction
In recent years, the increasing complexity of engineering systems and the growing performance demands in control applications have intensified the need for accurate nonlinear models. Discrete-time (DT) input-output (IO) models are one of the most commonly applied structures in system identification as they incorporate a broad spectrum of model classes, ranging from linear time-invariant (LTI) (Hespanha, 2018) and linear parameter-varying (LPV) (Tóth, 2010) models to various nonlinear structures (Schoukens and Ljung, 2019). First-principle (FP) models can be obtained in DT IO form based on known physical laws and engineering insight. While such physics-based models provide interpretable system descriptions, they often capture only the dominant dynamics, while additional high-complexity effects, such as friction properties or aerodynamic forces, are typically neglected.
@@
As an alternative approach, various data-driven identification methods have been developed for modeling nonlinear systems with an IO model structure (Schoukens and Ljung, 2019). In particular, recent advances employing deep artificial neural networks (ANNs) have demonstrated superior modeling accuracy compared to conventional approaches (Ljung et al., 2020). However, the practical use of ANN-based black-box models in control-oriented applications, e.g., trajectory planning, remains limited due to their lack of physical interpretability (Ljung, 2010). Furthermore, ANN-based models typically exhibit poor extrapolation capabilities beyond the range of the training data, and substantial learning effort is often spent on rediscovering system behaviors that are already well understood from first-principles knowledge.
@@
To address these challenges, different strategies have been proposed in the literature, starting from (light) grey-box modeling (Bohlin, 2006) till physics-informed neural networks (PINNs) (Raissi et al., 2019) and physics-guided neural networks (PGNNs) (Daw et al., 2022). One of the most promising directions of these hybrid approaches is model augmentation (Schön et al., 2022; Hoekstra et al., 2025). The augmentation approach aims at combining FP models, i.e., baseline models, with flexible learning components to achieve faster convergence and better model accuracy compared to black-box learning methods (Djeumou et al., 2022). Furthermore, model augmentation produces interpretable models with a clear understanding of how the learning component complements the baseline dynamics.
@@
In this paper, we investigate a widely used model augmentation structure in the literature, namely the additive formulation. This approach connects the physics-based and learning components in parallel and the corresponding parameters of the two components are jointly optimized. However, simultaneously tuning the learning-based and physical parameters results in the two ”subcomponents” competing with each other (Bolderman et al., 2022). As a result, the ANN can learn relations that could be represented by the FP model, while the baseline parameters can be tuned to physically unrealistic values. This effect undermines physical interpretability of the model estimate and can even compromise the extrapolation capabilities of the resulting model.
@@
This challenge was first addressed in Bolderman et al. (2022) in the context of PGNN-based feedforward control by introducing an additional regularization term into the cost function, penalizing deviations of the baseline parameters from their nominal values. This addition to the cost function effectively limits deviations of the baseline parameters compared to their initial values, hence preventing them from reaching a physically unrealistic parameter domain. Despite its simplicity, the approach has provided good experimental results (Bolderman et al., 2024); moreover, the method can be straightforwardly extended for more complex model augmentation structures (Hoekstra et al., 2025). Another attractive approach is based on an orthogonal projection-based regularization, introduced in Kon et al. (2022), also for feedforward control applications. This approach promotes a specific orthogonality between the baseline and learning components via regularization, penalizing when the ANN learns the already known relations represented by the baseline model. This technique has been adapted and generalized for nonlinear system identification in Györök et al. (2025); however, since these approaches promote orthogonality via regularization, inherently, there exists a trade-off between model accuracy and the desired complementarity. Finding the appropriate trade-off parameter (i.e., regularization coefficient) may not be intuitive. To overcome this, we propose a direct parametrization with guaranteed orthogonality between the baseline and learning components on a selected data set without requiring any trade-off parameter. We also show that under certain conditions for the dataset used to impose such an orthogonality property, the tuned baseline parameters converge to their physically true values.
@@ 2 Problem statement
where $f^{\mathrm{ANN}}$ here represents a fully connected, feedforward neural network with $\theta_{\mathrm{a}}\in\mathbb{R}^{n_{\theta_{\mathrm{a}}}}$ being the collection of its parameters. Alternatively, $f^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}$ can be replaced by any function approximator without loss of generality. Instead of the additive formulation, many other model augmentation structures can be selected from the literature, e.g., see Retzler et al. (2024). However, the additive, i.e., parallel, formulation offers a transparent model structure with clear separation between the baseline and learning components (Hoekstra et al., 2025), hence can be an attractive approach for practical applications.
To achieve the best possible data-fit, while simultaneously acquiring as accurate baseline parameters as possible, an efficient approach is to co-estimate $\theta_{\mathrm{b}}$ and $\theta_{\mathrm{a}}$ parameters, as proposed in Bolderman et al. (2022). Hence, using a data sequence $\mathcal{D}_{N}=\left\{(x_{i},y_{i})\right\}_{i=0}^{N-1}$ generated by (1), the parameters are estimated by minimizing the prediction error loss function expressed as
@@ 4.1 Recovery of the baseline parameters
If the structure of $\delta$ can be inferred from prior physical knowledge, it is possible to specifically design an input sequence that satisfies Condition 4 (at least with respect to the suspected unmodeled terms). This is a common scenario in practical applications. For example, in aerial robotics, the drag force is known to depend quadratically on the relative velocity of the object (Huang et al., 2024). However, the associated aerodynamic computations are often complex, and this force term is typically left unmodeled. Incorporating these physics-based insights can, however, guide the design of experiments that fulfil Condition 4.
@@ 4.2 Comparison with the standard parametrization
The error of the estimated baseline parameters in (25) is similar to the one in Donati et al. (2025), which also follows intuitively. When the dynamics of (14) are recovered on the training data, the exact baseline parameters are found only if the learning component identifies the unmodeled terms and nothing else. If the ANN learns parts that the baseline model could represent, the true baseline parameters can not be retained.
@@ 4.3 Consistency analysis
In the previous derivations, we assumed that the model training results in such parameters for which the exact relations of the data-generating system are recovered on the training set. Now, we will provide certain conditions under which this assumption is satisfied for $N\to\infty$, i.e., we will show consistency of the estimator following the arguments of Ljung (1978). For that, first we require that the data-generating system (1) is stable.
@@
The identification criterion given by (4) satisfies Condition C1 in Ljung (1978), hence the proof of (Ljung, 1978, Lemma 3.1) applies for the considered case. $\blacksquare$
@@
In practice, Condition 11 serves only as a technical condition. After acquiring the estimate of the model parameters $\hat{\theta}$, it can be numerically checked whether the given model realization satisfies this requirement. For a strict guarantee of satisfying Condition 11, a so-called stable-by-design parametrization can be applied. See, e.g., Revay et al. (2024); Kon et al. (2025). Extending these methods for the model augmentation setting is an open question and can be the topic of future research.
@@
For linear time-invariant models, the classical persistency of excitation (PE) condition that is characterized solely on the input trajectory in the open loop case, implies a fundamental informativity requirement, i.e., distinguishability of (equivalence sets of) parameters under the data (Gevers et al., 2009). For NARX-type models with linear-in-the-parameters form as (2), the PE condition requires the regressor matrix composed of nonlinear basis functions to be of full column rank. Unlike for linear systems, this condition depends not only on the input signal, but also on the resulting output trajectories. Consequently, PE requirements must be assessed at the level of the nonlinear regressor. Considering the full nonlinear model augmentation structure of (3) further complicates the PE condition. In practice, space-filling and model-based excitation designs (see, e.g., (Liu et al., 2025)) can be applied to select inputs for maximizing data information content over the regressor space.
@@
For the proof of (34), see Lemma 4.1 in Ljung (1978). Note that the applied loss function (4) fulfills Condition (4.4) in Ljung (1978). To prove (35), refer back to Theorem 7. With the conditions of Lemma 15, Assumption 6 is trivially satisfied; hence, (35) holds under Condition 4. $\blacksquare$
@@ 4.4 Covariance of the model parameters
Finally, we show that, due to the orthogonality between the baseline and learning components, the proposed parametrization results in zero covariance between the estimated $\theta_{\mathrm{b}}$ and $\theta_{\mathrm{a}}$ parameters. Under the conditions of the consistency results, the asymptotic distribution of $\hat{\theta}^{N}$ can be expressed w.r.t. to $\lim_{N\rightarrow\infty}\hat{\theta}^{N}=\theta^{\ast}\in\Theta^{\ast}$ as $\sqrt{N}(\hat{\theta}^{N}-\theta^{\ast})\in\mathrm{As}\ \mathcal{N}(0,P_{\theta})$ (Ljung, 1998). Then, under the considered quadratic loss function,
@@
where $\psi_{k}(\theta_{\ast})\in\mathbb{R}^{n_{\mathrm{y}}\times n_{\theta}}$ is the Jacobian matrix [^1] Note that $\psi_{k}$ is defined with a different dimensional notation compared to Ljung (1998) for practical reasons. $\partial\hat{y}_{k}/\partial\theta$. Based on $\mathcal{D}_{N}$ and parameter estimate $\hat{\theta}^{N}$, $P_{\theta}$ can be estimated as
@@ 5.1 Academic simulation example
where the baseline parameters are $\hat{\theta}_{1}$, and $\hat{\theta}_{2}$ with initial values as $\hat{\theta}_{1}^{0}=0.8$, $\hat{\theta}_{2}^{0}=0.03$. Similar physics-based baseline models have high relevance in system identification theory (Schoukens and Tiels, 2017). For the learning component, a simple feedforward ANN with 1 hidden layer and 16 neurons is applied, using the hyperbolic tangent (tanh) activation. The ANN parameters are initialized with the Xavier method (Glorot and Bengio, 2010).
Three distinct datasets are generated for training to highlight the importance of data generation, each with $N=1024$ data points. The first approach applies a white noise input signal with $u_{k}\sim\mathcal{N}(0,0.3^{2})$ for generating 512 samples, then the rest of the data points are acquired by using the first half of the input signal multiplied by -1. This results in a symmetric training data set (denoted by $\mathcal{D}_{N}^{\mathrm{(1)}}$), since the even nonlinearities are orthogonal to odd nonlinearities, according to (17), Condition 4 is satisfied. The second set $\mathcal{D}_{N}^{(2)}$ is gathered by applying 1024 input samples directly generated by the previous distribution. This results in a dataset with, most likely, small asymmetries. However, these asymmetries will disappear, and (17) holds for $N\to\infty$. Finally, $\mathcal{D}_{N}^{(3)}$ is acquired by using an asymmetric distribution of $u_{k}\sim\mathcal{N}(-0.01,0.3^{2})$, which will not satisfy Condition 4 when $N\to\infty$. For testing, a data set is constructed with $N_{\mathrm{test}}=1024$ using the same input distribution as in $\mathcal{D}_{N}^{(2)}$. For better demonstration of the results, the test data is kept noise-free. To minimize (4), we applied the Adam optimizer for 500 epochs, followed by the L-BFGS method for 1000 iterations, using the identification pipeline proposed in Bemporad (2025). For benchmarking, we applied the standard additive augmentation, as well as the proposed orthogonal-by-construction method, using 10 different initialization points.
@@ 5.2 Physics-motivated simulation example
Numerical parameter values corresponding to the data-generating system are $m=1$ kg, $k_{\mathrm{s}}=1000~\mathrm{N/m}$, $c_{\mathrm{d}}=100~\mathrm{Ns/m}$, $F_{\mathrm{C}}=0.8~\mathrm{N}$, $F_{\mathrm{S}}=1.2~\mathrm{N}$, $v_{\mathrm{S}}=0.02~\mathrm{m/s}$ with a sampling time of $T_{\mathrm{s}}=0.01$ s. These values result in $\theta_{1}\approx 0.48$, $\theta_{2}\approx-0.48$, $\theta_{3}=\theta_{4}\approx 4.8\cdot 10^{-5}$ (all in appropriate SI unit). Similar MSD interconnections have high importance in mechanical engineering practice, as they are used for dynamic modeling a wide range of components, e.g., suspension of a ground vehicle (Prabakar et al., 2013). These systems are usually described with linear models, which is convenient for system analysis and control design. Hence, we utilize the following physics-based baseline model:
@@
We apply the same ANN structure, initialization method, and training environment as in Sect. 5.1, but now with 2500 Adam and 500 L-BFGS epochs, respectively. For benchmarking, we apply the standard additive structure, but now complemented with the orthogonal regularization approach (see Kon et al. (2022); Györök et al. (2025)), which promotes a similar orthogonality as our method. Moreover, for further benchmarking, we apply another regularization-based technique (denoted as ”baseline regularization” in the remainder), adapted from Bolderman et al. (2022), which penalizes deviations of the baseline parameters compared to their nominal values. To demonstrate the advantage of our method over regularization-based techniques, we perform a Monte-Carlo study with 5 steps, using various trade-off parameter values (regularization coefficients). As the true baseline model parameters have different magnitudes, now we evaluate the estimation performance by the relative error metric, as $\|(\theta_{\mathrm{b}}^{\ast}-\hat{\theta}_{\mathrm{b}})/\theta_{\mathrm{b}}^{\ast}\|_{2}$. The results are shown in Fig. 5. As shown in the results, the orthogonal regularization–based approach converges toward the proposed method in terms of baseline model parameter estimation error as the trade-off parameter is increased. However, a slight reduction of the overall model performance can be noticed for the regularization-based approach. The baseline regularization method inherently depends on the initial baseline parameter values, since for large values of the trade-off parameter, $\hat{\theta}_{\mathrm{b}}$ remains close to the initial value $\hat{\theta}_{\mathrm{b}}^{0}$. Both approaches demonstrate that, as the regularization coefficient is reduced, the baseline parameters become non-unique and may deviate substantially from the true parameters $\theta_{\mathrm{b}}^{\ast}$. At the same time, lower trade-off parameter values can yield improved model performance, an effect that is particularly visible for the baseline regularization method. In practical system identification scenarios, where only test or validation errors are available during hyperparameter tuning, selecting appropriate regularization coefficients is therefore nontrivial and potentially a complex task. This observation highlights a key advantage of the proposed orthogonal-by-construction method, namely that it requires no hyperparameter tuning.